### v1

@@ 3.11 Detecting AGN through X-ray variability
Since X-ray emission is one of the most prominent AGN features, it has been pivotal for the discovery and study of the population of SMBH in the Universe [27]. This is because normal galaxies emit little radiation at these energies. Thus, the presence of X-ray emission, in excess of the one expected from the X-ray binaries and SNe in a galaxy, is among the most reliable methods to discover accreting SMBHs. In addition to the detection of significant X-ray flux from a distant source, variability can also play a significant role in its identification. This applies in particular to the elusive population of low-luminosity/low-accretion SMBH which are numerically more abundant than bright AGN, especially in the local universe [e.g. 9, 5, 293, 294].
The use of persistent variability to detect faint AGN has not been exploited in the X-ray band as much as in the optical band (Sect. 2.5). Nevertheless, it has been shown to be effective in pushing the limit of AGN detection in deep, multi-epoch X-ray surveys. In particular, the works of [320, 64] have exploited the repeated observations of the CDF-S spanning up to 17 years to identify LLAGN which escape classical selection criteria. As shown in Fig. 34, 10%-20% of sources with optical-to-X-ray flux ratio lower than 0.1 ( the lower limit usually adopted to confirm the presence of an AGN, Maccacaro et al. 188) are detected to be significantly variable. Such objects possess a luminosity erg/s and a (stacked) spectral power-law index typical of AGN. However, their BH mass and accretion rate are and times smaller than their bright counterparts, respectively. We point out that any deep multi-epoch survey has the potential of applying the same method to identify LLAGN if the survey cadence is properly planned. In fact, based on these examples, similar results can be expected from eROSITA or the future Athena mission.
@@
Swift/BAT provides continuous monitoring of many AGN but is not sufficiently sensitive to create densely sampled light curves, on short timescales, and in various energy bands. The available light curves can be used for a study of the average properties of the X-ray PSD at low frequencies. Analysis of the long term Swift/BAT light curves of the brightest AGN so far has confirmed that the slope of the power spectrum in AGN is below , and the PSD amplitude at energies above keV is similar to the PSD amplitude in the 2-10 keV band [272, 217]. However, the light curve errors are large even with a binning of one month, therefore it is not possible to investigate how the PSD properties (slopes, amplitude, ) depend on MBH and , as a function of energy.
The situation will improve with the Wide Field Monitor on board eXTP, which is a future satellite led by the Chinese Academy of Sciences and universities in China, with partnerships with various European institutions, and is planned to launch in 2027. The WFM will monitor hundreds of sources daily, but the number of AGN among these sources will not be very large. According to [328], the maximum sensitivity of the Wide Field Monitor will be mCrab in 50 ksec. If this is over the 2–50 keV energy range, this could imply a mCrab flux in the 2-10 keV band for power-law-like spectra with a slope of (the flux limit increases/decreases with steeper/flatter PL slopes). There are unobscured AGN in the nearby Universe with flux higher than mCrab [99]. So, WFM will help us to determine low-frequency PSD slopes, as well as amplitudes, for a sizable number of sources. We should be able to investigate their dependence on BH mass and/or accretion rate as well, although we will probably not be able to investigate PSD variations with energy and we will not be able to study the PSD bending frequency in detail. However, as before with data from other satellites, we will be able to use excess variance and VFP analysis (see Sect. 3.8) on sparsely sampled lightcurves of several hundred sources, especially in the regions of the sky observed repeatedly by the scanning strategy of the mission (see Fig. 14 of 117). This will give us the opportunity to investigate the variability properties of a much larger sample of objects, especially at higher redshifts [e.g. 159, 236].
@@
eROSITA (extended ROentgen Survey with an Imaging Telescope Array) is the primary instrument on-board the Russian-German ”Spectrum-Roentgen-Gamma” (SRG) mission. eROSITA was developed by the Max-Planck Institute for Extraterrestrial Physics (MPE) and was launched in 2019. Since December 2019, eROSITA has been conducting an all-sky survey in the X-ray band up to 10 keV, in which the entire celestial sphere will be mapped once every six months. Up to 2025, eROSITA has completed four of its eight intended all-sky survey passes. Consequently, eROSITA will provide light curves with a maximum of four points over two years for faint sources over most of the sky in the 0.2-8 keV band (rest frame). It is difficult to perform PSD analysis when there are few points in the light curves. Furthermore, the analysis of eROSITA data is intrinsically complex due to the changing PSF and sensitivity during the scans [22, 33]. However, significantly better sampled light curves will be available for sources near the poles, where eROSITA will observe AGN from tens to hundreds of times (see Fig.5.7.3 of 197 and updates on the mission web page141414https://erosita.mpe.mpg.de). Consequently, a detailed PSD analysis may be possible for these AGN. The big advantage of eROSITA when it comes to variability studies is the enormous number of AGN that will be detected during the sky survey. The very large number of AGN should allow for a detailed characterisation of X-ray variability in AGN through the use of statistics such as or any other statistic that could be an estimate of the integral of the PSD over the sampled timescales. eROSITA will benefit from redshift and BH mass estimates from several ancillary spectroscopic and photometric surveys, including LSST, which should allow us to better constrain the ensemble variability dependence on AGN properties, as demonstrated by the first published results [34, 87, 245].

### v3

@@ 3.11 Detecting AGN through X-ray variability
Since X-ray emission is one of the most prominent AGN features, it has been pivotal for the discovery and study of the population of SMBH in the Universe [28]. This is because normal galaxies emit little radiation at these energies. Thus, the presence of X-ray emission, in excess of the one expected from the X-ray binaries and SNe in a galaxy, is among the most reliable methods to discover accreting SMBHs. In addition to the detection of significant X-ray flux from a distant source, variability can also play a significant role in its identification. This applies in particular to the elusive population of low-luminosity/low-accretion SMBH which are numerically more abundant than bright AGN, especially in the local universe [e.g. 10, 5, 295, 296].
The use of persistent variability to detect faint AGN has not been exploited in the X-ray band as much as in the optical band (Sect. 2.5). Nevertheless, it has been shown to be effective in pushing the limit of AGN detection in deep, multi-epoch X-ray surveys. In particular, the works of [322, 65] have exploited the repeated observations of the CDF-S spanning up to 17 years to identify LLAGN which escape classical selection criteria. As shown in Fig. 34, 10%-20% of sources with optical-to-X-ray flux ratio lower than 0.1 ( the lower limit usually adopted to confirm the presence of an AGN, Maccacaro et al. 190) are detected to be significantly variable. Such objects possess a luminosity erg/s and a (stacked) spectral power-law index typical of AGN. However, their BH mass and accretion rate are and times smaller than their bright counterparts, respectively. We point out that any deep multi-epoch survey has the potential of applying the same method to identify LLAGN if the survey cadence is properly planned. In fact, based on these examples, similar results can be expected from eROSITA or the future Athena mission.
@@
Swift/BAT provides continuous monitoring of many AGN but is not sufficiently sensitive to create densely sampled light curves, on short timescales and in various energy bands. The available light curves can be used for a study of the average properties of the X-ray PSD at low frequencies. Analysis of the long term Swift/BAT light curves of the brightest AGN so far has confirmed that the slope of the power spectrum in AGN is below , and the PSD amplitude at energies above keV is similar to the PSD amplitude in the 2-10 keV band [274, 219]. However, the light curve errors are large even with a binning of one month, therefore it is not possible to investigate how the PSD properties (slopes, amplitude, ) depend on MBH and , as a function of energy.
The situation will improve with the Wide Field Monitor on board eXTP, which is a future satellite led by the Chinese Academy of Sciences and universities in China, with partnerships with various European institutions, and is planned to launch in 2027. The WFM will monitor hundreds of sources daily, but the number of AGN among these sources will not be very large. According to [330], the maximum sensitivity of the Wide Field Monitor will be mCrab in 50 ksec. If this is over the 2–50 keV energy range, this could imply a mCrab flux in the 2-10 keV band for power-law-like spectra with a slope of (the flux limit increases/decreases with steeper/flatter PL slopes). There are unobscured AGN in the nearby Universe with flux higher than mCrab [101]. So, WFM will help us to determine low-frequency PSD slopes, as well as amplitudes, for a sizable number of sources. We should be able to investigate their dependence on BH mass and/or accretion rate as well, although we will probably not be able to investigate PSD variations with energy and we will not be able to study the PSD bending frequency in detail. However, as before with data from other satellites, we will be able to use excess variance and VFP analysis (see Sect. 3.8) on sparsely sampled lightcurves of several hundred sources, especially in the regions of the sky observed repeatedly by the scanning strategy of the mission (see Fig. 14 of 119). This will give us the opportunity to investigate the variability properties of a much larger sample of objects, especially at higher redshifts [e.g. 161, 238].
@@
eROSITA (extended ROentgen Survey with an Imaging Telescope Array) is the primary instrument on-board the Russian-German ”Spectrum-Roentgen-Gamma” (SRG) mission. eROSITA was developed by the Max-Planck Institute for Extraterrestrial Physics (MPE) and was launched in 2019. Since December 2019, eROSITA has been conducting an all-sky survey in the X-ray band up to 10 keV, in which the entire celestial sphere will be mapped once every six months. Up to 2025, eROSITA has completed four of its eight intended all-sky survey passes. Consequently, eROSITA will provide light curves with a maximum of four points over two years for faint sources over most of the sky in the 0.2-8 keV band (rest frame). It is difficult to perform PSD analysis when there are few points in the light curves. Furthermore, the analysis of eROSITA data is intrinsically complex due to the changing PSF and sensitivity during the scans [23, 34]. However, significantly better sampled light curves will be available for sources near the poles, where eROSITA will observe AGN from tens to hundreds of times (see Fig.5.7.3 of 199 and updates on the mission web page141414https://erosita.mpe.mpg.de). Consequently, a detailed PSD analysis may be possible for these AGN. The big advantage of eROSITA when it comes to variability studies is the enormous number of AGN that will be detected during the sky survey. The very large number of AGN should allow for a detailed characterisation of X-ray variability in AGN through the use of statistics such as or any other statistic that could be an estimate of the integral of the PSD over the sampled timescales. eROSITA will benefit from redshift and BH mass estimates from several ancillary spectroscopic and photometric surveys, including LSST, which should allow us to better constrain the ensemble variability dependence on AGN properties, as demonstrated by the first published results [35, 89, 247].
@@ Acknowledgements
Declarations
Funding: I. Papadakis thanks the University of Naples Federico II for support through its “Visiting professor” program. M. Paolillo acknowledges support from the Italian PRIN - MIUR 2022 “SUNRISE” and the INAF grant TIMEDOMES. This research has been supported in part by the EU HORIZON MSCA Doctorate Network “TALES”.
Conflict of interest/Competing interests: the authors have no relevant financial or non-financial interests to disclose.
Ethics approval and consent to participate: not applicable.
Consent for publication: all the non-original material reproduced in this review is published under explicit consent of the original authors, journals or according to the individual journal reproduction permission policies.
Data availability: for access to the data cited in this review the reader is referred to the original publications.
Materials availability: not applicable.
Code availability: not applicable.
Author contribution: all authors have contributed equally to the preparation of this manuscript.